We are flipping a coin 6 times. What is the probability of the outcomes to be :

(Binomial distribution).

a) two tails.

b) At least 5 tails.

c) The most 2 heads


1
Expert's answer
2021-09-05T18:55:16-0400

Solution:

"p=q=\frac12"

"n=6"

"X\sim Bin(n,p)"

a) P(two tails)=P(X=2)"=^6C_2(0.5)^2(0.5)^4=0.234375"

b) At least 5 tails)="P(X\ge5)=P(X=5)+P(X=6)"

"=^6C_5(0.5)^5(0.5)^1+^6C_6(0.5)^6(0.5)^0
\\=6(0.5)^6+(0.5)^6
\\=0.109375"

c) P(The most 2 heads)"=P(X\le2)=P(X=0)+P(X=1)+P(X=2)"

"=^6C_0(0.5)^0(0.5)^6+^6C_1(0.5)^1(0.5)^5+^6C_2(0.5)^2(0.5)^4
\\=(0.5)^6+6(0.5)^6+15(0.5)^6
\\=0.34375"